OpenStudy (anonymous):

differentiable function of x; x^2+y^2-9=0 HELP!

OpenStudy (anonymous):

Not sure what you want. Take the square root of it all?

OpenStudy (anonymous):

Please explain what you need.....

OpenStudy (anonymous):

Do you mean to do implicit differentiation on that equation ????

OpenStudy (anonymous):

yes the implicit differentiation.

OpenStudy (anonymous):

\[\frac{d}{dx}(x^2+y^2) = \frac{d}{dx}(9)\]

OpenStudy (anonymous):

\[x^{2} + y^{2} - 9=0; \frac{ dy }{ dx }= \frac{ y-x }{ 3y-x }\]

OpenStudy (anonymous):

what's d/dx of x^2? what's d/dx of y^2? what's d/dx of 9?

OpenStudy (anonymous):

in respect to x

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

How about you answer my questions first.